A constant torque of 35 m∙N is applied to a wheel with a moment of inertia of 1.5 kg∙m2. (a) Find the angular acceleration.

(b) Find also the angular speed at the end of 25s, assuming the wheel starts from rest. Show your complete solution.


1
Expert's answer
2021-11-15T10:33:42-0500

(a) The angular acceleration is


"\alpha=\frac TI=23.3\text{ rad/s}."

(b) The angular speed is

"\omega=\frac12\alpha t^2=\frac12·23.3·25^2=7281\text{ rad/s}."